OpenStudy (zyi6):

Factor the Polynomial below: gx^2 - 3hx^2 - 6fy^2 -gy^2 + 6fx^2 + 3hy^2

OpenStudy (sshayer):

\[gx^2-3hx^2-6fy^2-gy^2+6fx^2+3hy^2=x^2(?-?+?)-y^2(?-?+?)\] \[=(?-?+?)(x^2-y^2)\] \[x^2-y^2=(x+y)(x-y)\] now complete the problem.